In a game called Taxation and Evasion, a player rolls a pair of dice. If on any turn the sum is 8, 10, or 12, the player gets audited. Otherwise, she avoids taxes. Suppose a player takes 5 turns at rolling the dice. The probability that she gets audited no more than 2 times is?

Please show all work (I must use Excel to find the answer)
Answer #1

P(getting 8,10 or 12 on pair of dice) =P(S =8)+P(S=10)+P(S=12) =(5/36)+(3/36)+(1/36)=9/36=1/4=0.25

here this is binomial with parameter n=5 and p=0.25

probability that she gets audited no more than 2 times is =P(X<=2) =binomdist(2,5,0.25,true)

=0.8965